Q.3 Tullahoma Company purchased equipment for $27,500. It depreciated the equipment over a five-year life by the double-declining-balance method until the end of the second year, at which time the asset was sold for $8,500. Calculate the gain or loss on the sale at the end of the second year.


 Q.4 Swanson & Hiller, Inc., purchased a new machine on September 1, 2008 at a cost of $108,000. The machine’s estimated useful life at the time of the purchase was five years, and its residual value was $8,000.

 

Instructions

a. Prepare a complete depreciation schedule, beginning with calendar year 2008, under each of

the methods listed below (assume that the half-year convention is used):

1. Straight-line.

2. 200 percent declining-balance.

3. 150 percent declining-balance, switching to straight-line when that maximizes the expense.




1
Expert's answer
2021-05-24T13:01:19-0400

QUESTION 3

Given;

Purchase cost of equipment = $27,500

Useful life = 5 years

Depreciation rate under double declining balance = (100 /5) x 2  

                                        = 40%

The depreciation schedule for the first and second year has been prepared as follows:






The calculation of loss on sale of equipment has been made as follows:

Loss on sale of equipment=Book value at the end of second year

−Sale value of asset

                                               =$9,900−$8,500

                                               =$1,400


Loss on sale of equipment is $1,400


QUESTION 4

information is given

Date of purchase = 1 September 2008

Cost of the machine = $108,000

Life of machine = 5 years

Residual value = $8,000

Statement showing schedule of depreciation by the straight-line method





Straight line method depreciation = Cost of the machine – Residual value

÷useful life

Straight line method depreciation = $108,000 − $8,000÷5

Straight line method depreciation = $20,000 Yearly

Straight line method depreciation (Half year) = $20,000÷2 = $10,000


Schedule of depreciation by 200 percent declining balance method

Information is given


Date of purchase = 1 September 2008

Cost of the machine = $108,000

Life of machine = 5 years

Residual value = $8,000


Rate of depreciation per year = 100÷5 years x 200% = 40%

Rate of depreciation for six month = 40% x ½ = 20%


Statement showing the schedule of depreciation by 200 percent declining-balance.





Schedule of depreciation by 150 percent declining balance switching to straight-line

Information is given

Date of purchase = 1 September 2008

Cost of the machine = $108,000

Life of machine = 5 years

Residual value = $8,000

Rate of depreciation = 100÷5 years x 150% 

Rate of depreciation = 30% per year

Rate of depreciation = 15% (For half year)

Statement showing the schedule of depreciation by 150 percent declining-balance.





1. Book value at the end of the year = Book value at the start of the year −

 Depreciation

2. . Switch to straight line method from year 5 , Depreciation expense calculated as (31,487 – 8,000 / 2) = $11,743 for both the remaining years